# Strassen Attention: Unlocking Compositional Abilities in Transformers Based on a New Lower Bound Method

### v2

@@ 3.2.2 The Match3 task
[…]
@@
[…]
@@
[…]
@@
$Sum_{2}[\ell,m](p)=$
$\begin{cases}1&\exists j,k\in[\ell]\text{ s.t. }p_{j}+p_{k}=-1\pmod{m},\\ 0&\text{otherwise}\end{cases}$
@@
For even $\ell$, it holds that:
[…]
@@
Indeed, first note that the value $m-2=2\ell-2$ is not used. Now, if $c_{i}=1$, then $p_{i}\oplus q$ has numbers $2i$ and $12\ell-2i-$ 1, summing up to $12\ell-$ 1. Next, if $c_{i}=0$, in $p_{i}\oplus q$ only two numbers can appear, $1$ and $2i\ell$, whose sum is neither $12\ell-$ 1 nor $14\ell-$ 1 because $i\leq\ell/2$. This completes the proof. ∎
@@ 3.2.3 Binary Relation Composition
More precisely, the output of the $(i,j)$-th token for $(i,j)\in[\sqrt{n}]^{2}$ has to be a real number $y_{ij}$ with $\textsf{sign}(y_{ij})=(B\circ A)_{ij}$. We refer to $A$ and $B$ as “relations” on the set $[\sqrt{n}]$, with $A_{ij}$ indicating whether the pair $(i,j)$ is in the relation $A$ and $B_{ij}$ doing so for relation $B$. For an example, imagine that $A=B$ encodes a relation for a group of researchers where two researchers $i$ and $j$ are related if they have a paper in co-authorship. Then the composition $B\circ A$ refers to the relation of “having a common co-author”.
@@ 4 Strassen attention – An efficient mechanism to solve complex tasks
we get an expression for $a_{i}$ in terms of their matrix products:
$a_{i}=\frac{(X\widehat{Y}Z)_{ii}}{(XYZ)_{ii}}.$
On top of exhibiting a faster implementation, we now show that, in contrast to standard attention, Strassen attention allows a one-layer Transformer to solve all the 3 tasks described in Section 3.2.
@@
The function composition and the binary relation composition tasks can be solved by a one-layer Transformer with 1 Strassen attention head, constant embedding dimension, and constant-size output MLP.
@@
Function composition. In the function composition task, we get a $(2n+1)$-length sequence of numbers
@@
$f_{i}=n\begin{pmatrix}(\phi(i))^{2}\\ 2\phi(i)\\ -1\\ 0\\ 0\\ 0\end{pmatrix},\qquad g_{j}=n\begin{pmatrix}-1\\ j\\ j^{2}\\ (\phi(j))^{2}\\ 2\phi(j)\\ -1\end{pmatrix},\qquad h_{k}=n\begin{pmatrix}0\\ 0\\ 0\\ -1\\ k-n\\ k^{2}-2kn+n^{2}\end{pmatrix}$
@@
In particular, the maximum of $a_{2n+1,j,k}$ is for $j$ and $k$ such that $j=\phi(2n+1)=x$, and $k=n+\phi(j)=n+\phi(x)=n+g(x)$, and other values of $a_{2n+1,j,k}$ are by an $e^{\Omega(n^{2})}$-factor smaller. Hence, with precision ±$\pm e^{-\Omega(n^{2})}$, we obtain $a_{2n+1}=v_{j}\odot v_{k}$ for $j=x$ and $k=n+g(x)$. Observe that $\phi(k)=\phi(n+g(x))=h(g(x))$, so it is enough for $v_{j}\odot v_{k}$ to have a coordinate equal to $\phi(k)$. We can achieve this by setting matrices $V_{1},V_{2}$ in (17) such that the first coordinates of $v_{j}$ and $v_{k}$ are 1 and $\phi(k)$, respectively.
Binary relation composition. In this task, the length of input is a square number $n=m^{2}$, and tokens are indexed by pairs $(i,j)\in[m]^{2}$. The token, indexed by $(i,j)$, receives on input two bits $A_{ij},B_{ij}$ from two Boolean matrices $A,B\in\{0,1\}^{m\times m}$. As an output, the $(i,j)$-th token has to produce a real number $y_{ij}$ such that
@@
The expression in brackets has the “integral part”, and also has the term $\frac{cm^{3}+dm^{2}+km+\ell}{m^{5}}$ which is $O(1/m)$ and is different for different quadruples $(c,d,k,\ell)$. Hence, there is a unique quadruple $(c,d,k,\ell)$, establishing the maximum of the above expression, and it also maximizes the “integral part” of (25), i.e., $\left[-(i-c)^{2}-(d-k)^{2}-(\ell-j)^{2}+A_{cd}+B_{k\ell}\right]$. Because of the factor $n^{4}$, the value for the other quadruples will be smaller by at least $\Omega(n^{4}/m^{5})=\Omega(n^{3/2})$.
@@
[…]
@@
Since all numbers under consideration are polynomial in $n$, the output $a_{i}$ will be equal to $v_{j}\odot v_{k}$ for the maximal pair $(j,k)$ up to $\exp\{-\Omega(n)\}$-precision. In the output MLP, we have to check if $p_{i}+p_{j}+p_{k}=\Sigma$ for this pair $(j,k)$. It is enough to define $V_{1}$, $V_{2}$ so that the 5th and the 6th coordinate of $v_{j}$ and $v_{k}$ are $1,p_{j}$ and $p_{k},1$, respectively. As a result, the 2nd, the 5h, and the 6th coordinates of $x_{i}+a_{i}$ will be $-p_{i}$, $p_{k}$, and $p_{j}$, respectively, allowing us to find out if $p_{i}+p_{j}+p_{k}=\Sigma$ with a constant-size output MLP. ∎
@@ 5 Disentangling Strassen from Standard and Triangular attentions
So far, we have evaluated tasks that are challenging for one-layer standard attention Transformers but (in principle) easy for one-layer Strassen attention Transformers. In this section, we extend our analysis to triangular attention. As a reminder, running the triangular attention mechanism on a general sequence of length $n$, requires the creation of $n^{2}$ tokens. In this regime, the triangular attention running time becomes $n^{3}$ However, when the input is already structured as a $\sqrt{n}\times\sqrt{n}$ matrix, one can run the triangular attention on it directly, making the running time $O(n^{3/2})$. One such task example using structured input is the binary relation composition task. In this case, a one-layer triangular attention can perform this task with one attention head, constant embedding dimension and constant-size output MLP.
We devise a variant of the binary relation task that allows us to disentangle the performance of Strassen attention with that of the triangular and standard attentions, namely the quotient binary relation composition task. The latter takes as inputs two Boolean matrices $A,B\in\{0,1\}^{m\times m}$ and a “coloring” function $\textsf{col}\colon[m]\to[m]$, where $m=\sqrt{n}$. There are $n=m^{2}$ input tokens, indexed by pairs from $[m]^{2}$, with the $(i,j)$-th token receiving $A_{i,j},B_{i,j}$ and $(\textsf{col}(i),\textsf{col}(j))$ as inputs. The quotient of the composition $B\circ A$ by col is a Boolean matrix $B\circ A/\textsf{col}\in\{0,1\}^{m\times m}$, defined by:
@@
The quotient binary relation composition task is solvable by a one-layer Transformer with 1 Strassen-attention head, constant embedding dimension and constant-size output MLP. At the same time, this task cannot be solved by any one-layer Transformer $T$ with $n^{o(1)}$ standard-attention heads, $n^{o(1)}$ triangular-attention heads, $n^{o(1)}$ embedding dimension, and $n^{o(1)}$-size output MLP.
@@
Upper bound. The upper bound is very similar to our construction for the binary relation composition task in Theorem 4.2. Namely, first we replace $-(d-k)^{2}$ by $-(\textsf{col}(d)-\textsf{col}(k))^{2}$ in (25). After this modification, the maximum of (25) will be attained on a single quadruple $(c,d,k,\ell)$, and for this quadruple we will have $c=i,\ell=j$, $\textsf{col}(d)=\textsf{col}(k)$ and $A_{id}=B_{kj}=1$ if and only if a quadruple, satisfying these equalities, exists. However, due to the definition of our task, we have to add a smaller term, enforcing that among quadruples, satisfying this property, those with $d\neq k$ have a larger value of (25). This can be achieved by adding a term of the form $(d-k)^{2}/m^{3}$, which is always $O(1/m)$ so that the largest possible difference in this term is smaller than the smallest possible difference of the “integral part” in (25).
@@
Lower bound. We employ the same technique as in Theorem 3.2. However, we cannot rely on it directly as now we have to deal with the triangular attention.
@@
for $j\in[s]$, and letting all the other entries of the matrices $A,B$ to be 00. The coloring function is defined by $\textsf{col}(1)=\textsf{col}(2)=1$ and
@@
[…]
Figure 2: An example of the construction. Nodes apart from 1 and 2 are split into 2 equal groups – 3, 4, 5, 6 and 7, 8, 9, 10. If $A_{ij}=1$ (resp., $B_{ij}=1$), we draw an $A$-labeled (resp., a $B$-labeled) edge from $i$ to $j$. The $A$-edges can only go from 1 to 3,4,5,6, and the word $p$ determines, which of these edges are present. Likewise, the $B$-edges only go from 7, 8, 9, 10 to 2, according to whether $q$ has 1 or 0 in the corresponding position. Nodes 3 and 7, 4 and 8, 5 and 9, 6 and 10 have the same color but different pairs have a different color. Therefore, the only way we can have $(B\circ A/c)_{12}=1$ is when 1 has an $A$-edge to some node on the left, and the node of the same color from the right has a $B$-edge to 2. In the example from the figure, this is true for 5 and 9 (which happens because $p$ and $q$ both have 1 in the third position).
@@
Then we define a hypothesis class $H$ by considering parts of (28) that depend on $p$ as inputs and parts that depend on $q$ as parameters. On the one hand, since $d,H$ and the size of $\mathcal{N}$ are assumed to be $n^{o(1)}$, the VC dimension of this class is $n^{o(1)}$ by Theorem 2.3 in [9]. On the other hand, its VC dimension is lower bounded by the VC dimension of the set of columns of the matrix $\mathrm{Disj}(p,q)$, which is at least $s=n/2-1$ as established in the proof of Proposition 3.9. ∎
@@ 6 Experiments and Results
In this section, we systematically compare the performance of standard, triangular, third-order and Strassen attention in four tasks: (i) Indicator of the 1st coordinate in the Function Composition (with $f=g$ ), (ii) Binary Relation Composition (with $A=B$), (iii) Match3 (over position-aware permutations) and (iv) Quotient Binary Relation Composition (with $B=A^{T}$). To obtain a tighter upper bound on the capability of the standard attention, we chose to implement simple special cases of the tasks (see Appendix A for all the details on data generation). To compare these attention mechanisms, we evaluate their performance (accuracy per epoch) and training time (accuracy per minute) on the test sets of these tasks. Furthermore, Appendix B provides a thorough analysis in terms of computational performance, evaluating forward pass times, GPU and memory usage. Code for our experiments can be found at https://anonymous.4open.science/r/StrassenAttention-70D8/.
@@
7 Future directions and limitations
Our results pave the way for several directions of future research. First, we have introduced a new method to obtain lower bounds for one-layer Transformer, based on splitting the VC dimension. We hope that the method could be applied to other architectures and other tasks involving complex reasoning in more naturalistic settings (e.g., semantic parsing). Second, given the differences observed in learning stability between the third-order and Strassen attentions, the latter seems to be associated with a smoother loss landscape, an hypothesis that needs to be confirmed and studied. Third, Strassen attention can be adapted to incorporate interactions that involve more than three tokens, possibly capturing more complex patterns in data. Yet, practical learnability of such higher-order interactions needs to be assessed. Finally, and related to the previous point, although the main goal of our work was to gain a deeper theoretical understanding of the abilities of the Transformer, our conclusions are limited by using toy tasks. Our next step is to test the theoretical advantages of the Strassen attention using specific benchmarks or even in learning methods such as masked language modeling. At the same time, the possible applications of the Strassen attention are not limited to compositionality but could extend to such areas as knowledge graphs, protein structure prediction and others.
Appendix A Experimental Setup
A.1 Datasets

### v3

@@ 3.2.2 The Match3 task
Next, we define the Match[$_{3}[n,m$] task [20]. It is a sequence-to-sequence task. The input is presented to the Transformer as a sequence of $n$ tokens, encoding a vector of integers $(p_{1},\ldots,p_{n})\in[m-1]^{n}$. The output is a vector $(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}$, required to satisfy:
@@
Let $T$ be any one-layer standard-attention Transformer with $size(T)=n^{o(1)}$. Then $T$ cannot solve the Match[$_{3}[n,m$] task for m=2n2m=2n-2.
@@
We fix the value of $p_{1}$ to be equal to 1, and consider the output of Match[$_{3}[n,m$] at the first position. If we additionally set $p_{j}\neq m-2$ for $j\geq 2$, we obtain $p_{1}+p_{j}+p_{k}=0\pmod{m}$ if and only if $j,k\geq 2$ and $p_{j}+p_{k}=-1\pmod{m}$. Hence, a Transformer for the Match[$_{3}[n,m$] task can be converted into a Transformer, computing the following function in the auxiliary token:
@@
$Sum_{2}[\ell,m](p):=$ $\begin{cases}1&\exists j,k\in[\ell]\text{ s.t. }p_{j}+p_{k}=-1\pmod{m},\\ 0&\text{otherwise}\end{cases}$
@@
For even $\ell$, it holds that $\textsf{split-VC}(Sum_{2}[\ell,2\ell])\geq\ell/2.$
@@
Indeed, first note that the value $m-2=\ell-2$ is not used. Now, if $c_{i}=1$, then $p_{i}\oplus q$ has numbers $2i$ and $2\ell-2i-1$, summing up to $2\ell-1$. Next, if $c_{i}=0$, in $p_{i}\oplus q$ only two numbers can appear, $1$ and $2i\ell$, whose sum is neither $2\ell-1$ nor $4\ell-1$ because $i\leq\ell/2$. This completes the proof. ∎
@@ 3.2.3 Binary Relation Composition
More precisely, the output of the $(i,j)$-th token for $(i,j)\in[\sqrt{n}]^{2}$ has to be a real number $y_{ij}$ with $\textsf{sign}(y_{ij})=(B\circ A)_{ij}$. We refer to $A$ and $B$ as “relations” on the set $\sqrt{n}]$, with $A_{ij}$ indicating whether the pair $(i,j)$ is in the relation $A$ and $B_{ij}$ doing so for relation $B$. For an example, imagine that $A=B$ encodes a relation for a group of researchers where two researchers $i$ and $j$ are related if they have a paper in co-authorship. Then the composition $B\circ A$ refers to the relation of “having a common co-author”.
@@ 4 Strassen attention – An efficient mechanism to solve complex tasks
we get an expression for $a_{i}$ in terms of their matrix products $a_{i}=\frac{(X\widehat{Y}Z)_{ii}}{(XYZ)_{ii}}$. ∎
On top of exhibiting a faster runnin-time, we now show that, in contrast to standard attention, Strassen attention allows a one-layer Transformer to solve all the 3 tasks described in Section 3.2.
@@
We consider these three tasks separately.
Function composition
In the function composition task, we get a $(2n+1)$-length sequence of numbers
@@
$f_{i}=n\begin{pmatrix}(\phi(i))^{2}\\ 2\phi(i)\\ -1\\ 0\\ 0\\ 0\end{pmatrix},\qquad g_{j}=n\begin{pmatrix}-1\\ j\\ j^{2}\\ (\phi(j))^{2}\\ 2\phi(j)\\ -1\end{pmatrix}$,
$h_{k}=n\begin{pmatrix}0\\ 0\\ 0\\ -1\\ k-n\\ k^{2}-2kn+n^{2}\end{pmatrix}$
@@
In particular, the maximum of $a_{2n+1,j,k}$ is for $j$ and $k$ such that $j=\phi(2n+1)=x$, and $k=n+\phi(j)=n+\phi(x)=n+g(x)$, and other values of $a_{2n+1,j,k}$ are by an $e^{\Omega(n^{2})}$-factor smaller. Hence, with precision $\pm e^{-\Omega(n^{2})}$, we obtain $a_{2n+1}=v_{j}\odot v_{k}$ for $j=x$ and $k=n+g(x)$. Observe that $\phi(k)=\phi(n+g(x))=h(g(x))$, so it is enough for $v_{j}\odot v_{k}$ to have a coordinate equal to $\phi(k)$. We can achieve this by setting matrices $V_{1},V_{2}$ in (15) such that the first coordinates of $v_{j}$ and $v_{k}$ are 1 and $\phi(k)$, respectively.
Binary relation composition
In this task, the length of input is a square number $n=m^{2}$, and tokens are indexed by pairs $(i,j)\in[m]^{2}$. The token, indexed by $(i,j)$, receives on input two bits $A_{ij},B_{ij}$ from two Boolean matrices $A,B\in\{0,1\}^{m\times m}$. As an output, the $(i,j)$-th token has to produce a real number $y_{ij}$ such that
@@
The expression in brackets has the “integral part”, and also has the term $\frac{cm^{3}+dm^{2}+km+\ell}{m^{5}}$ which is $O(1/m)$ and is different for different quadruples $(c,d,k,\ell)$. Hence, there is a unique quadruple $(c,d,k,\ell)$, establishing the maximum of the above expression, and it also maximizes the “integral part” $\left[-(i-c)^{2}-(d-k)^{2}-(\ell-j)^{2}+A_{cd}+B_{k\ell}\right]$. Because of the factor $n^{4}$, the value for the other quadruples will be smaller by at least $\Omega(n^{4}/m^{5})=\Omega(n^{3/2})$.
@@
Match3
On input, we get an array $p_{1}\ldots p_{n}\in[m-1]^{n}$. We first describe how to check, for a fixed $\Sigma$ and for every $i=1,\ldots,n$, if there exist $j,k\in[n]$ such that $p_{i}+p_{j}+p_{k}=\Sigma$ using one Strassen attention head. We get a solution for the Match[$_{3}[n,m$] task with 2 attention heads by applying this construction to $\Sigma=m$ and $\Sigma=2m$.
@@
Since all numbers under consideration are polynomial in $n$, the output $a_{i}$ will be equal to $v_{j}\odot v_{k}$ for the maximal pair $(j,k)$ up to $\exp\{-\Omega(n)\}$-precision. In the output MLP, we have to check if $p_{i}+p_{j}+p_{k}=\Sigma$ for this pair $(j,k)$. It is enough to define $V_{1}$, $V_{2}$ so that the 5th and the 6th coordinate of $v_{j}$ and $v_{k}$ are $1,p_{j}$ and $p_{k},1$, respectively. As a result, the 2nd, the 5h, and the 6th coordinates of $x_{i}+a_{i}$ will be $p_{i}$, $p_{k}$, and $p_{j}$, respectively, allowing us to find out if $p_{i}+p_{j}+p_{k}=\Sigma$ with a constant-size output MLP. ∎
@@ 5 Disentangling Strassen from Standard and Triangular attentions
So far, we have evaluated tasks that are challenging for one-layer standard attention Transformers but (in principle) easy for one-layer Strassen attention Transformers. In this section, we extend our analysis to triangular attention. As a reminder, running the triangular attention mechanism on a general sequence of length $n$, requires the creation of $n^{2}$ tokens. In this regime, the triangular attention running time becomes $n^{3}$ . However, when the input is already structured as a $\sqrt{n}\times\sqrt{n}$ matrix, one can run the triangular attention on it directly, making the running time $O(n^{3/2})$. One such task example using structured input is the binary relation composition task. In this case, a one-layer triangular attention can perform this task with one attention head, constant embedding dimension and constant-size output MLP.
We devise a variant of the binary relation task that allows us to disentangle the performance of Strassen attention with that of the triangular and standard attentions, namely the quotient binary relation composition task. The latter takes as inputs two Boolean matrices $A,B\in\{0,1\}^{m\times m}$ and a “coloring” function $\textsf{col}\colon[m]\to[m]$, where $m=\sqrt{n}$. There are $n=m^{2}$ input tokens, indexed by pairs from $m]^{2}$, with the $(i,j)$-th token receiving $A_{i,j},B_{i,j}$ and $(\textsf{col}(i),\textsf{col}(j))$ as inputs. The quotient of the composition $B\circ A$ by col is a Boolean matrix $B\circ A/\textsf{col}\in\{0,1\}^{m\times m}$, defined by:
@@
The quotient binary relation composition task is solvable by a one-layer Strassen-attention constant-size transformer. At the same time, this task cannot be solved by any one-layer Transformer $T$ with $n^{o(1)}$ standard-attention heads, $n^{o(1)}$ triangular-attention heads, $n^{o(1)}$ embedding dimension, and $n^{o(1)}$-size output MLP.
@@
We start with the upper bound.
Upper bound
The upper bound is very similar to our construction for the binary relation composition task in Theorem 4.2. Namely, first we replace $(d-k)^{2}$ by $(\textsf{col}(d)-\textsf{col}(k))^{2}$ in (23). After this modification, the maximum of (23) will be attained on a single quadruple $(c,d,k,\ell)$, and for this quadruple we will have $c=i,\ell=j$, $\textsf{col}(d)=\textsf{col}(k)$ and $A_{id}=B_{kj}=1$ if and only if a quadruple, satisfying these equalities, exists. However, due to the definition of our task, we have to add a smaller term, enforcing that among quadruples, satisfying this property, those with $d\neq k$ have a larger value of (23). This can be achieved by adding a term of the form $(d-k)^{2}/m^{3}$, which is always $O(1/m)$ so that the largest possible difference in this term is smaller than the smallest possible difference of the “integral part” in (23).
@@
Lower bound
We employ the same technique as in Theorem 3.2. However, we cannot rely on it directly as now we have to deal with the triangular attention.
@@
for $j\in[s]$, and letting all the other entries of the matrices $A,B$ to be 0. The coloring function is defined by $\textsf{col}(1)=\textsf{col}(2)=1$ and
@@
1122443355668877991010AAAABBBBBBp=1010p=1010q=0111q=0111
Figure 2: An example of the construction. Nodes apart from 1 and 2 are split into 2 equal groups – 3, 4, 5, 6 and 7, 8, 9, 10. If $A_{ij}=1$ (resp., $B_{ij}=1$), we draw an AA-labeled (resp., a $B$-labeled) edge from $i$ to $j$. The AA-edges can only go from 1 to 3,4,5,6, and the word $p$ determines, which of these edges are present. Likewise, the $B$-edges only go from 7, 8, 9, 10 to 2, according to whether $q$ has 1 or 0 in the corresponding position. Nodes 3 and 7, 4 and 8, 5 and 9, 6 and 10 have the same color but different pairs have a different color. Therefore, the only way we can have $(B\circ A/c)_{12}=1$ is when 1 has an AA-edge to some node on the left, and the node of the same color from the right has a $B$-edge to 2. In the example from the figure, this is true for 5 and 9 (which happens because $p$ and $q$ both have 1 in the third position).
@@
Then we define a hypothesis class $H$ by considering parts of (26) that depend on $p$ as inputs and parts that depend on $q$ as parameters. On the one hand, since $d,H$ and the size of $\mathcal{N}$ are assumed to be $n^{o(1)}$, the VC dimension of this class is $n^{o(1)}$ by Theorem 2.3 in [9]. On the other hand, its VC dimension is lower bounded by the VC dimension of the set of columns of the matrix $\mathrm{Disj}(p,q)$, which is at least s=n/21s=n/2-1 as established in the proof of Proposition 3.9. ∎
@@ 6 Experiments and Results
In this section, we systematically compare the performance of standard, triangular, third-order and Strassen attention in four tasks: (i) Indicator of the 1st coordinate in the Function Composition (with $f=g$ ), (ii) Binary Relation Composition (with $A=B$), (iii) Match3 (over position-aware permutations) and (iv) Quotient Binary Relation Composition (with $B=A^{T}$). To obtain a tighter upper bound on the capability of the standard attention, we chose to implement simple special cases of the tasks (see Section 7 for all the details on data generation). To compare these attention mechanisms, we evaluate their performance (accuracy per epoch) and training time (accuracy per minute) on the test sets of these tasks. Furthermore, Section 8 provides a thorough analysis in terms of computational performance, evaluating forward pass times, GPU and memory usage. Code for our experiments can be found at [Uncaptioned image] strassen-attention-neurips25-434F/.
@@
7 Experimental Setup
7.1 Datasets